### 10 Mar 2026 v1

# YOLO-NAS-Bench: A Surrogate Benchmark with Self-Evolving Predictors for YOLO Architecture Search

Abstract
Neural Architecture Search (NAS) for object detection is severely bottlenecked by high evaluation cost, as fully training each candidate YOLO architecture on COCO demands days of GPU time. Meanwhile, existing NAS benchmarks largely target image classification, leaving the detection community without a comparable benchmark for NAS evaluation. To address this gap, we introduce YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. YOLO-NAS-Bench defines a search space spanning channel width, block depth, and operator type across both backbone and neck, covering the core modules of YOLOv8 through YOLO12. We sample 1,000 architectures via random, stratified, and Latin Hypercube strategies, train them on COCO-mini, and build a LightGBM surrogate predictor. To sharpen the predictor in the high-performance regime most relevant to NAS, we propose a Self-Evolving Mechanism that progressively aligns the predictor’s training distribution with the high-performance frontier, by using the predictor itself to discover and evaluate informative architectures in each iteration. This method grows the pool to 1,500 architectures and raises the ensemble predictor’s $R^{2}$ from 0.770 to 0.815 and Sparse Kendall Tau from 0.694 to 0.752, demonstrating strong predictive accuracy and ranking consistency. Using the final predictor as the fitness function for evolutionary search, we discover architectures that surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s discriminative power for top-performing detection architectures.
1 Introduction
Figure 1: Latency vs. mAP on COCO-mini. Architectures discovered by our predictor-guided EA search consistently Pareto-dominate all official YOLO baselines (v8–v12) across the full latency spectrum, demonstrating the strong discriminative power of the YOLO-NAS-Bench surrogate predictor.
Neural Architecture Search (NAS) demonstrates notable success in automating the design of high-performing image classifiers [28, 17, 19]. However, extending NAS to object detection remains expensive: detection models are much larger, datasets such as COCO [16] require far more compute per training run, and the complete search space needs to jointly consider backbone, neck, and head components. Fully training a single YOLO-family architecture on COCO can take days on a multi-GPU cluster. Evaluating thousands of candidates—as NAS algorithms typically demand—thus becomes infeasible.
NAS benchmarks significantly accelerate research in the classification domain by providing precomputed look-up tables or surrogate predictors, thereby decoupling NAS algorithm development from the expensive architecture evaluation loop and offering a unified benchmark for fair comparison. NAS-Bench-101 [24] and NAS-Bench-201 [6] offer tabular benchmarks with exhaustive evaluation of small cell-based search spaces. NAS-Bench-301 [25] scales to the full DARTS space via a surrogate predictor. Yet all these benchmarks target image classification—their search spaces, training pipelines, and evaluation protocols cannot be directly transferred to detection architectures. Meanwhile, detection-specific NAS methods such as Det-NAS [5], OPANAS [15], and YOLO-NAS [1] each define bespoke search spaces and evaluation setups, making fair cross-method comparison difficult. The community still lacks unified benchmarks for object detection NAS.
We address this gap with YOLO-NAS-Bench, the first surrogate benchmark designed specifically for YOLO-style object detectors. YOLO-NAS-Bench features three tightly integrated components: (1) a comprehensive search space spanning channel width, block depth, and operator type across both backbone and neck, covering the core building blocks of YOLOv8 through YOLO12; (2) an initial ground-truth database of 1,000 architectures sampled via complementary strategies and fully trained on COCO-mini under a unified protocol; and (3) a Self-Evolving Predictor that iteratively enriches the training pool with high-value architectures discovered by evolutionary search, culminating in an ensemble of 10 LightGBM models that achieves sKT = 0.752 and $R^{2}$ = 0.815. We validate the predictor’s practical utility by using its predicted mAP as the fitness function for EA-based search [19]. As shown in Fig. 1, the discovered architectures surpass all official YOLO baselines (v8–v12) at comparable latency.
Our contributions are summarized as follows:
  • We design a YOLO-oriented search space covering backbone and neck with channel, depth, and operator dimensions that span the key modules of YOLOv8–YOLO12. Through random, stratified, and Latin Hypercube sampling, we build a benchmark database of 1,000 architectures fully trained on COCO-mini.
  • We propose a Self-Evolving Predictor that bridges the distribution gap between uniformly sampled training data and the high-performance frontier critical to NAS. The predictor guides latency-bucketed evolutionary search to discover promising architectures, which are trained and fed back to retrain the predictor, forming a self-evolving loop. Over 10 rounds, the pool grows to 1,500 architectures and sKT rises from 0.694 to 0.752.
  • We demonstrate the predictor’s practical utility by deploying it as the fitness function for EA search and discovering architectures that surpass official YOLO baselines (v8–v12) at comparable latency on COCO-mini.
2 Related Work
Real-Time Object Detectors.
The YOLO family has driven rapid progress in real-time detection. YOLOv8 [11] introduces the C2f module with streamlined cross-stage feature fusion. YOLOv9 [23] proposes GELAN and Programmable Gradient Information (PGI) for improved gradient flow. YOLOv10 [22] eliminates the NMS post-processing step via consistent dual assignments. YOLO11 [12] advances efficiency with C3k2 re-parameterized blocks and C2PSA attention modules. YOLO12 [20] adopts an attention-centric design, pushing the accuracy–speed Pareto frontier further. Beyond the YOLO lineage, RT-DETR [27] demonstrates that Transformer-based detectors can also achieve real-time speeds. Despite these continuous improvements, architecture design in the detection domain remains predominantly manual, motivating the need for automated search and, consequently, a standardized benchmark to evaluate NAS algorithms fairly.
NAS Benchmarks.
Tabular benchmarks such as NAS-Bench-101 [24] (423 k architectures on CIFAR-10) and NAS-Bench-201 [6] (15,625 cell-based architectures across three datasets) are instrumental in democratizing NAS research for image classification by removing the evaluation bottleneck. NAS-Bench-301 [25] scales to the full DARTS search space by training a surrogate predictor evaluated via the Sparse Kendall Tau (sKT) metric, establishing a protocol that subsequent benchmarks adopt. NAS-Bench-360 [21] broadens the scope to diverse tasks and data modalities, yet does not include detection-specific search spaces. YOLOBench [14] characterizes a fixed set of YOLO model families but cannot be adapted for NAS algorithm evaluation. Existing benchmarks predominantly target classification or evaluate fixed model families. Consequently, surrogate benchmarks for detection-oriented NAS remain scarce, and our work addresses this gap.
NAS for Object Detection.
Several works apply NAS to detection sub-problems: NAS-FPN [9] searches for feature pyramid topologies; Det-NAS [5] performs backbone architecture search with a one-shot supernet; OPANAS [15] optimizes FPN neck architectures via an accuracy predictor; SP-NAS [10] proposes a serial-to-parallel search strategy for detection backbones; and YOLO-NAS [1] employs the proprietary AutoNAC framework to design YOLO architectures. Each of these methods defines its own search space, training recipe, and evaluation setup, making direct cross-method comparison difficult. Our paper addresses this fragmentation by providing a shared search space, a precomputed architecture–performance database, and a calibrated surrogate predictor that any NAS algorithm can query at near-zero cost.
3 Method
Figure 2: Overview of the YOLO-NAS-Bench pipeline. (1) A YOLO-style search space spanning channel, depth, and operator dimensions across both backbone and neck is defined. (2) 1,000 architectures are sampled via three complementary strategies and trained on COCO-mini. (3) A LightGBM predictor is trained on the resulting {architecture, mAP} pairs. (4) The Self-Evolving Predictor iteratively expands the pool with high-value architectures discovered by evolutionary search, and retrains the predictor over 10 rounds, yielding an ensemble of 10 LightGBM models over 1,500 architectures.
An overview of the YOLO-NAS-Bench is illustrated in Fig. 2. Starting from a carefully designed search space over YOLO-style architectures, we sample and fully train 1,000 architectures on COCO-mini to build a ground-truth performance database. Each configuration is converted into a compact feature vector, upon which a LightGBM surrogate predictor is trained. To further strengthen the predictor in the high-performance regime that matters most for NAS, we propose a Self-Evolving Predictor loop that iteratively discovers, trains, and assimilates promising architectures.
3.1 Benchmark Construction
Search Space Design.
We design a comprehensive search space that covers both the backbone and the neck of YOLO-style detectors, while keeping the detection head fixed. The space is parameterized along three dimensions:
  • Channel width: Each of the four backbone stages (P2–P5) has an independently selectable channel count. The candidate sets grow with the stage depth to reflect the natural widening of feature hierarchies in modern detectors.
  • Block depth: The number of repeated blocks within each stage is searchable, controlling the representation capacity at each resolution level.
  • Operator type: This dimension subsumes both the feature-extraction module in each stage and the downsampling operator between stages. Feature-extraction candidates include C2f [11], C3k2 and C2PSA [12], and C2fCIB [22]—the core building blocks employed across YOLOv8 through YOLO12—spanning a spectrum from lightweight convolutions (C2f) to efficient re-parameterized blocks (C3k2, C2fCIB) and attention-augmented modules (C2PSA). Downsampling candidates include standard Conv and the stride-channel-decoupled SCDown [22].
    In the backbone, the operator palette widens progressively: P2/P3 stages choose between {C2f, C3k2}, P4 adds C2fCIB, and P5 further includes C2PSA, reflecting the increasing complexity budget at lower resolutions. In the neck, channel width and depth are fixed to avoid excessive growth of the search space; only the operator type is searchable. Feature blocks N1–N4 share a single global choice from {C2f, C3k2, C2fCIB}, and the two downsample blocks D1–D2 share a single choice from {Conv, SCDown}. Upsample layers use fixed nn.Upsample throughout the YOLO family and are not searchable.
Tab. 1 shows the full search space specification. The combinatorial product of all dimensions yields an architecture space on the order of millions of unique configurations, providing a rich landscape for NAS algorithm evaluation.
Table 1: Search space of YOLO-NAS-Bench. Channel, depth, and operator choices for each stage.
Dimension Stage Candidates #
Channel P2 128, 192, 256 3
P3 256, 320, 384, 448, 512 5
P4111P4 channels must be at least P3 to avoid degenerate configurations. 384, 512, 640, 768 4
P5 768, 1024, 1280 3
Depth P2 1, 2 2
P3 2, 3, 4 3
P4 2, 3, 4 3
P5 2, 3 2
Operator (backbone) P2/P3 C2f, C3k2 2
P4 C2f, C3k2, C2fCIB 3
P5 C2f, C3k2, C2fCIB, C2PSA 4
Operator (neck) N1–N4 C2f, C3k2, C2fCIB 3
D1–D2 Conv, SCDown 2
Architecture Sampling.
To build a diverse and representative ground-truth database, we employ three complementary sampling strategies over the search space: (i) Random sampling (200 architectures) provides a uniform baseline coverage; (ii) Stratified sampling (400 architectures) bins candidates by Parameters and draws uniformly within each stratum, ensuring balanced representation across model scales; (iii) Latin Hypercube Sampling [18] (400 architectures) maximizes coverage across the high-dimensional discrete space by applying stratified sampling independently along each search dimension. The three strategies are complementary: random sampling covers the space broadly, stratified sampling prevents under-representation of light or heavy architectures, and LHS provides near-optimal space-filling properties.
All 1,000 sampled architectures are trained from scratch on COCO-mini under an identical training protocol (detailed in Sec. 4.1). COCO-mini is a class- and size-stratified 10% subset of COCO that preserves the original category and bounding-box size distributions. The resulting database constitutes the foundation of YOLO-NAS-Bench.
Architecture Encoding and Surrogate Predictor.
Each architecture configuration is converted into a 24-dimensional feature vector. Channel widths and block depths are represented as scalar values, while operator choices are one-hot encoded, allowing the tree-based predictor to split on each operator independently without imposing artificial ordering.
[…]
  • Coefficient of determination ($R^{2}$): measures overall regression quality.
  • Sparse Kendall Tau (sKT): the Kendall $\tau$ rank correlation computed after rounding predictions to $0.1\%$ precision, which discounts ranking changes due to negligible prediction noise. Formally, $\mathrm{sKT}=\tau\!\bigl(\mathbf{y},\,\lfloor\hat{\mathbf{y}}\rceil_{0.001}\bigr)$.
Figure 3: Self-Evolving Predictor. Starting from 1,000 architectures, the loop partitions latency into 10 buckets. For each bucket, EA search selects the top 5 architectures using predicted mAP as fitness and real latency as constraint. In each round, these 50 new architectures are trained on COCO-mini, merged into the pool, and the predictor is retrained. After 10 rounds the pool grows to 1,500 architectures, yielding an enhanced benchmark enriched in the high-performance regime most relevant to NAS.
3.2 Self-Evolving Predictor
A predictor trained on uniformly sampled architectures may under-represent the high-performance frontier, where ranking accuracy is most critical for NAS. To address this distribution mismatch, we propose a Self-Evolving Mechanism. As shown in Fig. 3, it iteratively enriches the training pool with architectures that the current predictor considers promising, thereby sharpening its discrimination in the region that matters.
Latency bucketing. The latency range observed across all 1,000 initial architectures is evenly partitioned into 10 buckets. This stratification ensures that the self-evolving loop discovers high-performing architectures at every latency operating point, not only at a single scale.
Evolutionary expansion (single round). Within each latency bucket, a target latency is sampled uniformly. An evolutionary algorithm (EA) [19] is then run with predicted mAP as the fitness function and real measured latency as the constraint. The EA uses a population of 50 and runs for 100 generations, adopts top 25% as parents, with offspring generated by uniform crossover (50%) and mutation (50%, mutation probability 0.2), and preserves the top 10% as elite. The final top 5 architectures from each bucket are retained on COCO-mini, yielding 50 new candidate architectures per round.
Iterative refinement. Each batch of 50 new architectures is fully trained on COCO-mini under the same protocol, their configuration and ground-truth mAP are recorded, and the pairs are merged into the architecture pool. The predictor is then retrained on the expanded pool. This loop repeats for 10 rounds, growing the pool from 1,000 to 1,500 architectures. Crucially, each round’s newly added architectures are biased toward the high-performance regions identified by the current predictor, progressively reducing the gap between the predictor’s training distribution and the distribution of architectures encountered during actual NAS search.
Ensemble prediction. After the final round, 10 LightGBM models are trained on the full 1,500-architecture pool with different random seeds. The ensemble prediction is the arithmetic mean of the 10 models, which reduces variance and further stabilizes ranking quality.
4 Experiments
4.1 Experimental Setup
Dataset. We construct COCO-mini by stratified sampling 10% of COCO [16] images, preserving the original category distribution and bounding-box size ratios. The resulting subset retains 80 classes and follows the standard train2017 / val2017 split. All architectures—both the benchmark pool and YOLO baselines—are trained and evaluated on this identical dataset, ensuring fair comparison.
[…]
Table 2: Unified training configuration for all architectures in YOLO-NAS-Bench.
Parameter Value
Epochs 120
Batch size 128
Image size $640\times 640$
Learning rate ($\text{lr}_{0}$) 0.01
LR schedule Step (decay at epoch 100)
Mosaic 1.0
MixUp 0.15
Copy-Paste 0.5
Pretrained No
Workers 16
Latency measurement. All latencies are measured on a single NVIDIA P40 GPU with batch size 1, $640{\times}640$ FP32 input, 10 warmup and 50 timed forward passes. The mean inference time is reported in milliseconds.
Predictor training. The surrogate predictor is trained with RMSE as the loss function. We report $R^{2}$ and sKT on a held-out 20% validation split.
4.2 Main Results
Predictor Quality.
Tab. 3 summarizes the surrogate predictor’s accuracy before and after Self-Evolving. The 10-model LightGBM ensemble trained on the initial 1,000 architectures achieves a validation sKT of 0.694. After 10 rounds of Self-Evolving, both $R^{2}$ and sKT get improved: $R^{2}$ rises from 0.770 to 0.815 (+4.5%), and sKT from 0.694 to 0.752 (+5.8%). An sKT of 0.752 indicates strong ranking consistency, while the $R^{2}$ of 0.815 suggests that the predictor captures the majority of variance in architecture performance. These results show that our surrogate predictor is a high-fidelity proxy for the true performance landscape, and confirm that the Self-Evolving mechanism effectively sharpens the predictor’s ranking ability. Fig. 4 further visualizes the close agreement between predicted and actual mAP across all 1,500 architectures.
Figure 4: Predicted vs. ground-truth mAP on the full 1,500-architecture pool. Each point is an architecture colored by its sampling source. Points cluster closely around the y=xy{=}x diagonal, confirming strong agreement between the ensemble predictor and ground-truth performance.
Table 3: Predictor quality before and after Self-Evolving. Metrics are reported on the validation split (20%). Ensemble: 10 homogeneous LightGBM models.
Setting #Archs $\boldsymbol{R^{2}}$ sKT
Before Self-Evolving 1,000 0.770 0.694
After Self-Evolving 1,500 0.815 0.752
Predictor-Guided Search Results.
Table 4: Predictor-guided EA search results vs. official YOLO baselines on COCO-mini. All models are trained from scratch under the same protocol. Latency is measured on a single P40 GPU. The top block shows architectures discovered by our predictor-guided EA; subsequent blocks show official baselines.
Large Medium-Large Medium Small
Model mAP (%) Lat. (ms) mAP (%) Lat. (ms) mAP (%) Lat. (ms) mAP (%) Lat. (ms)
[…]
YOLO12 32.9 54.13 30.8 42.51 30.4 22.17 27.2 20.40
YOLO11 32.8 37.73 32.1 29.84 30.8 20.03 27.7 15.30
YOLOv10 30.5 36.84 30.1 25.83 29.7 20.01 25.8 15.89
YOLOv9 32.3 53.07 31.6 26.85 31.6 25.59 27.1 19.25
YOLOv8 32.9 33.71 30.8 23.30 29.4 17.21 26.2 10.95
To validate that our surrogate predictor can reliably identify top-performing architectures, we conduct an evolutionary architecture search (EA) [19] using the ensemble predictor’s predicted mAP as the fitness function and real measured latency as the constraint. The EA searches within the same search space defined in Sec. 3.1. Top candidate architectures are then fully retrained on COCO-mini under the identical protocol of Tab. 2, producing ground-truth mAP values.
As shown in Tab. 4, the four predictor-discovered architectures Pareto-dominate all official YOLO baselines (v8–v12) across the full latency spectrum, achieving consistently higher mAP at equal or lower latency. The advantage is most pronounced in the small-model regime, where Arch-D surpasses YOLO11s by +4.2% mAP at comparable latency, and remains substantial at the large end, where Arch-A exceeds YOLO12x in mAP while being $1.5{\times}$ faster. These results confirm that the predictor not only achieves high ranking correlation (sKT) on held-out data, but also exhibits strong discriminative power in the top-performance regime—architectures it deems promising indeed yield superior performance after full retraining.
Table 5: Predictor type comparison on initial 1,000 architectures (no ensemble, same data split).
Predictor $\boldsymbol{R^{2}}$ sKT
LightGBM 0.768 0.699
XGBoost 0.758 0.696
NGBoost 0.755 0.704
Random Forest 0.744 0.678
MLP 0.053 0.440
4.3 Ablation Studies
Predictor type comparison.
Tab. 5 compares five predictor types trained on the initial 1,000-architecture pool: LightGBM [13], XGBoost [4], NGBoost [7], Random Forest [3], and MLP. Among the tree-based methods, LightGBM achieves the best balance of $R^{2}$ (0.768) and sKT (0.699), followed closely by NGBoost (sKT 0.704) and XGBoost ($R^{2}$ 0.758). Random Forest lags slightly behind. The MLP baseline performs poorly. These results justify our choice of LightGBM as the base model for the predictor.
Self-Evolving vs. random expansion.
To disentangle whether the predictor improvement stems from the Self-Evolving mechanism or merely from increasing the training pool size, we conduct an ablation: we train 200 randomly sampled architectures on COCO-mini, add them to the initial 1,000 architectures, and retrain the ensemble predictor. For fair comparison, we also truncate Self-Evolving when the pool reaches 1,200 architectures. Tab. 6 reports the results. Random expansion yields $R^{2}=0.776$ and sKT=0.701=0.701, while Self-Evolving achieves $R^{2}=0.798$ and sKT=0.738=0.738. The clear gap (+0.022 $R^{2}$, +0.037 sKT) confirms that the gain is attributable to the targeted enrichment of high-performance architectures by the Self-Evolving loop, rather than to the increase in data size alone.
Table 6: Self-Evolving vs. random pool expansion. Same pool size (1,200); metrics on validation split (20%).
Expansion strategy #Archs $\boldsymbol{R^{2}}$ sKT
Initial (no expansion) 1,000 0.770 0.694
Random +200 1,200 0.776 0.701
Self-Evolving +200 1,200 0.798 0.738
4.4 Limitations and future work.
The current benchmark is built on COCO-mini (10% of COCO) and measures latency on a single GPU (NVIDIA P40). To ensure the authenticity of latency measurements across different platforms, we do not model a latency predictor in the same way as mAP; thus, when using this benchmark for NAS, one still needs to measure the latency of each architecture empirically. But this cost is acceptable compared to the error that a latency predictor would introduce on unseen hardware. Extending YOLO-NAS-Bench to the full COCO dataset, diverse hardware platforms (edge GPUs, mobile NPUs), and additional tasks such as instance segmentation and pose estimation are natural next steps.
5 Conclusion
We present YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. By designing a search space that spans channel width, block depth, and operator type across backbone and neck, and by training 1,000 architectures sampled by different strategies on COCO-mini, we establish an initial performance database. Based on this, our Self-Evolving Predictor iteratively enriches the database with high-value architectures, raising the ensemble LightGBM predictor’s sKT from 0.694 to 0.752. Architectures discovered by predictor-guided evolutionary search surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s practical utility.

### 18 May 2026 v2

@@
 Using the final predictor as the fitness function for evolutionary search, we discover architectures that surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s discriminative power for top-performing detection architectures.
+The code is available at https://github.com/VDIGPKU/YOLO-NAS-Bench.
 1 Introduction
@@
   • We design a YOLO-oriented search space covering backbone and neck with channel, depth, and operator dimensions that span the key modules of YOLOv8–YOLO12.
-Through random, stratified, and Latin Hypercube sampling, we build a benchmark database of 1,000 architectures fully trained on COCO-mini.
+Through random, stratified, and Latin Hypercube sampling, we build an initial database of 1,000 architectures fully trained on COCO-mini.
   • We propose a Self-Evolving Predictor that bridges the distribution gap between uniformly sampled training data and the high-performance frontier critical to NAS.
@@
 3 Method
-Figure 2: Overview of the YOLO-NAS-Bench pipeline.
+Figure 2: Overview of YOLO-NAS-Bench.
 (1) A YOLO-style search space spanning channel, depth, and operator dimensions across both backbone and neck is defined.
@@
 (4) The Self-Evolving Predictor iteratively expands the pool with high-value architectures discovered by evolutionary search, and retrains the predictor over 10 rounds, yielding an ensemble of 10 LightGBM models over 1,500 architectures.
-An overview of the YOLO-NAS-Bench is illustrated in Fig.
+An overview of YOLO-NAS-Bench is illustrated in Fig.
 2.
@@
 We design a comprehensive search space that covers both the backbone and the neck of YOLO-style detectors, while keeping the detection head fixed. The space is parameterized along three dimensions:
-  • Channel width: Each of the four backbone stages (P2–P5) has an independently selectable channel count.
+  • Channel width: Each of the four backbone stages (S2–S5) has an independently selectable channel count.
 The candidate sets grow with the stage depth to reflect the natural widening of feature hierarchies in modern detectors.
@@
   • Operator type: This dimension subsumes both the feature-extraction module in each stage and the downsampling operator between stages.
-Feature-extraction candidates include C2f [11], C3k2 and C2PSA [12], and C2fCIB [22]—the core building blocks employed across YOLOv8 through YOLO12—spanning a spectrum from lightweight convolutions (C2f) to efficient re-parameterized blocks (C3k2, C2fCIB) and attention-augmented modules (C2PSA).
+Feature-extraction candidates include C2f [11], C3k2 [12], C2fCIB [22], and A2C2f [20]—the core building blocks employed across YOLOv8 through YOLO12—spanning a spectrum from lightweight convolutions (C2f) to efficient re-parameterized blocks (C3k2, C2fCIB) and attention-augmented modules (A2C2f).
 Downsampling candidates include standard Conv and the stride-channel-decoupled SCDown [22].
-    In the backbone, the operator palette widens progressively: P2/P3 stages choose between {C2f, C3k2}, P4 adds C2fCIB, and P5 further includes C2PSA, reflecting the increasing complexity budget at lower resolutions.
+In the backbone, the operator palette widens progressively: S2/S3 stages choose between {C2f, C3k2}, S4 adds C2fCIB, and S5 further includes A2C2f, reflecting the increasing complexity budget at lower resolutions.
 In the neck, channel width and depth are fixed to avoid excessive growth of the search space; only the operator type is searchable.
-Feature blocks N1–N4 share a single global choice from {C2f, C3k2, C2fCIB}, and the two downsample blocks D1–D2 share a single choice from {Conv, SCDown}.
+Feature blocks N1–N4 share a single global choice from {C2f, C3k2, C2fCIB, A2C2f}, and the two downsample blocks D1–D2 share a single choice from {Conv, SCDown}.
 Upsample layers use fixed nn.Upsample throughout the YOLO family and are not searchable.
@@
 Dimension Stage Candidates #
-Channel P2 128, 192, 256 3
-P3 256, 320, 384, 448, 512 5
-P4111P4 channels must be at least P3 to avoid degenerate configurations. 384, 512, 640, 768 4
-P5 768, 1024, 1280 3
-Depth P2 1, 2 2
-P3 2, 3, 4 3
-P4 2, 3, 4 3
-P5 2, 3 2
-Operator (backbone) P2/P3 C2f, C3k2 2
-P4 C2f, C3k2, C2fCIB 3
-P5 C2f, C3k2, C2fCIB, C2PSA 4
-Operator (neck) N1–N4 C2f, C3k2, C2fCIB 3
+Channel S2 128, 192, 256 3
+S3 256, 320, 384, 448, 512 5
+S4111S4 channels must be at least S3 to avoid degenerate configurations. 384, 512, 640, 768 4
+S5 768, 1024, 1280 3
+Depth S2 1, 2 2
+S3 2, 3, 4 3
+S4 2, 3, 4 3
+S5 2, 3 2
+Operator (backbone) S2/S3 C2f, C3k2 2
+S4 C2f, C3k2, C2fCIB 3
+S5 C2f, C3k2, C2fCIB, A2C2f 4
+Operator (neck) N1–N4 C2f, C3k2, C2fCIB, A2C2f 4
 D1–D2 Conv, SCDown 2
@@
 Architecture Encoding and Surrogate Predictor.
-Each architecture configuration is converted into a 24-dimensional feature vector.
+Each architecture configuration is converted into a 25-dimensional feature vector.
 Channel widths and block depths are represented as scalar values, while operator choices are one-hot encoded, allowing the tree-based predictor to split on each operator independently without imposing artificial ordering.
@@
 An evolutionary algorithm (EA) [19] is then run with predicted mAP as the fitness function and real measured latency as the constraint.
-The EA uses a population of 50 and runs for 100 generations, adopts top 25% as parents, with offspring generated by uniform crossover (50%) and mutation (50%, mutation probability 0.2), and preserves the top 10% as elite.
-The final top 5 architectures from each bucket are retained on COCO-mini, yielding 50 new candidate architectures per round.
+The EA runs for 100 generations with a population of 50, selecting the top 25% as parents, generating offspring via crossover (50%) or mutation (50%, rate 0.2), and carrying forward the top 10% as elite.
+The final top 5 architectures from each bucket are selected as candidates, yielding 50 new architectures per round.
 Iterative refinement.
@@
 This loop repeats for 10 rounds, growing the pool from 1,000 to 1,500 architectures.
-Crucially, each round’s newly added architectures are biased toward the high-performance regions identified by the current predictor, progressively reducing the gap between the predictor’s training distribution and the distribution of architectures encountered during actual NAS search.
+Crucially, each round’s newly added architectures are biased toward high-performance regions identified by the current predictor, progressively narrowing the distribution gap between predictor training and actual NAS search.
 Ensemble prediction.
@@
 4.1 Experimental Setup
-Dataset.
-We construct COCO-mini by stratified sampling 10% of COCO [16] images, preserving the original category distribution and bounding-box size ratios.
-The resulting subset retains 80 classes and follows the standard train2017 / val2017 split.
-All architectures—both the benchmark pool and YOLO baselines—are trained and evaluated on this identical dataset, ensuring fair comparison.
-[…]
 Table 2: Unified training configuration for all architectures in YOLO-NAS-Bench.
@@
 Workers 16
+Dataset.
+We construct COCO-mini by stratified sampling 10% of COCO [16] images, preserving the original category distribution and bounding-box size ratios.
+The resulting subset retains 80 classes and follows the standard train2017/val2017 split.
+All architectures—both the benchmark pool and YOLO baselines—are trained and evaluated on this identical dataset, ensuring fair comparison.
+[…]
 Latency measurement.
@@
 The 10-model LightGBM ensemble trained on the initial 1,000 architectures achieves a validation sKT of 0.694.
-After 10 rounds of Self-Evolving, both $R^{2}$ and sKT get improved: $R^{2}$ rises from 0.770 to 0.815 (+4.5%), and sKT from 0.694 to 0.752 (+5.8%).
+After 10 rounds of Self-Evolving, both $R^{2}$ and sKT improve: $R^{2}$ rises from 0.770 to 0.815 (+4.5%), and sKT from 0.694 to 0.752 (+5.8%).
 An sKT of 0.752 indicates strong ranking consistency, while the $R^{2}$ of 0.815 suggests that the predictor captures the majority of variance in architecture performance.
@@
 After Self-Evolving 1,500 0.815 0.752
-Predictor-Guided Search Results.
 Table 4: Predictor-guided EA search results vs.
@@
 The top block shows architectures discovered by our predictor-guided EA; subsequent blocks show official baselines.
-Large Medium-Large Medium Small
+Small (S) Medium (M) Large (L) Extra-Large (X)
 Model mAP (%) Lat. (ms) mAP (%) Lat. (ms) mAP (%) Lat. (ms) mAP (%) Lat. (ms)
 […]
-YOLO12 32.9 54.13 30.8 42.51 30.4 22.17 27.2 20.40
-YOLO11 32.8 37.73 32.1 29.84 30.8 20.03 27.7 15.30
-YOLOv10 30.5 36.84 30.1 25.83 29.7 20.01 25.8 15.89
-YOLOv9 32.3 53.07 31.6 26.85 31.6 25.59 27.1 19.25
-YOLOv8 32.9 33.71 30.8 23.30 29.4 17.21 26.2 10.95
+YOLO12 27.2 20.40 30.4 22.17 30.8 42.51 32.9 54.13
+YOLO11 27.7 15.30 30.8 20.03 32.1 29.84 32.8 37.73
+YOLOv10 25.8 15.89 29.7 20.01 30.1 25.83 30.5 36.84
+YOLOv9 27.1 19.25 31.6 25.59 31.6 26.85 32.3 53.07
+YOLOv8 26.2 10.95 29.4 17.21 30.8 23.30 32.9 33.71
+Predictor-Guided Search Results.
 To validate that our surrogate predictor can reliably identify top-performing architectures, we conduct an evolutionary architecture search (EA) [19] using the ensemble predictor’s predicted mAP as the fitness function and real measured latency as the constraint.
@@
 4, the four predictor-discovered architectures Pareto-dominate all official YOLO baselines (v8–v12) across the full latency spectrum, achieving consistently higher mAP at equal or lower latency.
-The advantage is most pronounced in the small-model regime, where Arch-D surpasses YOLO11s by +4.2% mAP at comparable latency, and remains substantial at the large end, where Arch-A exceeds YOLO12x in mAP while being $1.5{\times}$ faster.
+The advantage is most pronounced in the small-model regime, where Ours-s surpasses YOLO11s by +4.2% mAP at comparable latency, and remains substantial at the large end, where Ours-x exceeds YOLO12x in mAP while being $1.5{\times}$ faster.
 These results confirm that the predictor not only achieves high ranking correlation (sKT) on held-out data, but also exhibits strong discriminative power in the top-performance regime—architectures it deems promising indeed yield superior performance after full retraining.
@@
 For fair comparison, we also truncate Self-Evolving when the pool reaches 1,200 architectures.
-Tab.
-6 reports the results.
-Random expansion yields $R^{2}=0.776$ and sKT=0.701=0.701, while Self-Evolving achieves $R^{2}=0.798$ and sKT=0.738=0.738.
+As shown in Tab.
+6, random expansion yields $R^{2}=0.776$ and sKT=0.701=0.701, while Self-Evolving achieves $R^{2}=0.798$ and sKT=0.738=0.738.
 The clear gap (+0.022 $R^{2}$, +0.037 sKT) confirms that the gain is attributable to the targeted enrichment of high-performance architectures by the Self-Evolving loop, rather than to the increase in data size alone.
@@
 Self-Evolving +200 1,200 0.798 0.738
-4.4 Limitations and future work.
+4.4 Limitations and Future Work
 The current benchmark is built on COCO-mini (10% of COCO) and measures latency on a single GPU (NVIDIA P40).
-To ensure the authenticity of latency measurements across different platforms, we do not model a latency predictor in the same way as mAP; thus, when using this benchmark for NAS, one still needs to measure the latency of each architecture empirically.
+To ensure the authenticity of latency measurements across different platforms, we do not build a surrogate latency predictor as we do for mAP; thus, when using this benchmark for NAS, one still needs to measure the latency of each architecture empirically.
 But this cost is acceptable compared to the error that a latency predictor would introduce on unseen hardware.
@@
 We present YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors.
-By designing a search space that spans channel width, block depth, and operator type across backbone and neck, and by training 1,000 architectures sampled by different strategies on COCO-mini, we establish an initial performance database.
-Based on this, our Self-Evolving Predictor iteratively enriches the database with high-value architectures, raising the ensemble LightGBM predictor’s sKT from 0.694 to 0.752.
+By designing a search space that spans channel width, block depth, and operator type across backbone and neck, and by training 1,000 architectures sampled via complementary strategies on COCO-mini, we establish an initial performance database.
+Building on this, our Self-Evolving Predictor iteratively enriches the database with high-value architectures, raising the ensemble LightGBM predictor’s sKT from 0.694 to 0.752.
 Architectures discovered by predictor-guided evolutionary search surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s practical utility.
+6 Acknowledgments
+This work was supported by National Natural Science Foundation of China under Grant 62176007.